the quantitative subtest of the gre has a mean of 500 and a standard deviation of 100. what quantitative gre scores are so extreme that we expect to observe them in 1% or less of all test takers
Answer #1

Solution :

Using standard normal table ,

P(Z z) = 1%.

P(Z -2.33) = 0.01

z = -2.33

Using z-score formula,

x = z * +

x = -2.33 * 100 + 500 = 267

gre scores = 267